OpenStudy (anonymous):

>>>use elimination to solve system of equations<<< x+8y=3 4x-2y=7

jimthompson5910 (jim_thompson5910):

notice how -2y can be multiplied by 4 to get -8y, which means 8y + (-8y) =0y = 0, so that goes away so if you multiply everything in the second row by 4 you get this new equation 16x - 8y = 28 and you'll have this new system x+8y=3 16x-8y=28 now add the equations x+8y=3 16x-8y=28 ----------- 17x+0y=31 17x = 31 solve this for x. Then use this to find the solution for y

OpenStudy (anonymous):

(31/7,41/68)??

jimthompson5910 (jim_thompson5910):

x should be 31/17 not 31/7

jimthompson5910 (jim_thompson5910):

and y is incorrect, so try it again

OpenStudy (anonymous):

Idk? I got the same thing

jimthompson5910 (jim_thompson5910):

pick any equation with x and y in it, plug in x = 31/17 and solve for y x+8y=3 31/17+8y=3 ... plug in x = 31/17 8y = 3 - 31/17 8y = 51/17 - 31/17 8y = 20/17 y = (20/17)*(1/8) y = 20/136 y = 5/34 So x = 31/17 and y = 5/34

OpenStudy (anonymous):

So you can plug it in to any equation and still gt th same answer?

jimthompson5910 (jim_thompson5910):

yes because both equations were used to solve for x, which means that both x and y should make both equations true at the same time

jimthompson5910 (jim_thompson5910):

so that's why you can pick any equation

OpenStudy (anonymous):

Okay I know how to do them I just can't do them...

jimthompson5910 (jim_thompson5910):

just keep practicing and you'll get better